Determine if n=161304001 is a strong pseudo prime with respect to the base 2.


1
Expert's answer
2021-02-24T07:36:57-0500

Let "n=161304001"

"n-1=161304000=2^6\cdot2520375\\
2^{2520375}\equiv 30275048 \not\equiv 1 \text{ mod }161304001\\
2^{2 \cdot 2520375}\equiv -1 \text
{ mod } 161304001"

This shows that n is composite and also passes the Miller-Rabi test. Hence, n=161304001 is a strong pseudoprime.